• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name LeRose

    Log entry time 16:00:51 on May14,2004

    Entry number 121714

    keyword=Shift Summary

    Very quiet, no beam trips observed, no beam either!
    Hall C did their bleed through studies all day. We had no beam.
    At shift's end everything is ready to go.

    Bob Michaels calls at 15:53 to say that the Sun machines are being rebooted (after installation of security patches). Follow instructions on the sticky notes to start them up again.

    15:55 Jack is out of the Hall.

    15:56 MCC informed we are ready to go to "power permit".